# A Survey of Reinforcement Learning For Economics

### 09 Mar 2026 v1

@@ 6 $\sim 10^{8}$ Failed
The results are sobering for RL proponents. In settings where classical solutions exist and can be computed, RL adds little value and often underperforms. The base-stock policy, derived analytically for idealized settings, remains highly competitive when properly calibrated. RL struggles particularly with multi-echelon coupling, where orders at upstream stages affect costs at downstream stages many periods later, creating difficult credit assignment problems. [^88] Credit assignment refers to the difficulty of determining which past actions caused a delayed reward or cost. In multi-echelon systems, an upstream ordering decision today may not affect customer-facing costs for many periods, making it difficult for RL to learn the causal connection. The authors report that DRL required millions of training transitions to approach performance levels that base-stock policies achieve through closed-form calculation. The value proposition for RL in inventory management lies in problems where analytical solutions are unavailable, such as non-stationary demand, complex operational constraints, or settings where the cost structure does not admit tractable decomposition.
@@ 6.6 Real-Time Bidding
Real-time bidding for display advertising presents a budget pacing problem in which an advertiser must allocate a fixed budget across a sequence of auctions to maximize conversions. Wu et al. (2018) formalized this as an MDP where the state includes remaining budget, time elapsed, and recent win rates, with actions being bid multipliers. The RL agent learns to adapt bidding intensity to market conditions, increasing bids when competition is low and conserving budget when prices spike. Simulation experiments calibrated to real RTB data showed improvements over rule-based pacing algorithms.
@@ 6.7 Simulation Study: Bus Engine Replacement
I extend the single-engine problem to a fleet of $N$ engines with a capacity constraint limiting replacements per period. The state $s=(m_{1},\ldots,m_{N})$ records discretized mileage for each engine. Actions are subsets of engines to replace, subject to the capacity constraint. The per-period cost is $c(s,a)=\alpha\sum_{i}m_{i}+\beta|a|$, where $\alpha$ is the operating cost per unit mileage and $\beta$ is the replacement cost. [^89] The mileage-dependent operating cost $\alpha\sum_{i}m_{i}$ follows Rust’s original specification $c(x,\theta_{1})=\theta_{11}x$, creating a non-trivial threshold replacement policy. The fleet extension uses deterministic mileage increments to isolate the combinatorial scaling challenge that arises from the joint state of multiple engines. Mileage evolves deterministically; replaced engines reset to $m=0$; others increment by one bin. With $M=6$ mileage bins, the state space is $6^{N}$, yielding 1,296 states at $N=4$, 7,776 at $N=5$, 46,656 at $N=6$. Value iteration is feasible for $N\leq 5$.
@@ 7 Structural Estimation with Reinforcement Learning
Several recent papers have used RL training loops, namely Q-learning, temporal-difference learning, policy gradient, and actor-critic methods, to solve structural economic models at scales where conventional dynamic programming fails. [^90] I exclude papers that use neural network function approximation without an RL training mechanism. Inverse reinforcement learning is treated in the sister survey (Rust and Rawat, 2026). Throughout this chapter I adopt a unified notation. An MDP is a tuple $(\mathcal{S},\mathcal{A},P,r,\gamma)$ where $\mathcal{S}$ is the state space, $\mathcal{A}$ is the action space, $P(s^{\prime}|s,a)$ is the transition kernel, $r(s,a)$ is the per-period reward, and $\gamma\in[0,1)$ is the discount factor. [^91] Several of the papers reviewed here use $\beta$ for the discount factor, following economics convention. I translate all results to $\gamma$ for consistency with the RL literature and the rest of this survey. A policy $\pi:\mathcal{S}\to\Delta(\mathcal{A})$ maps states to distributions over actions. The value function under $\pi$ is $V^{\pi}(s)=\mathbb{E}_{\pi}\left[\sum_{t=0}^{\infty}\gamma^{t}r(s_{t},a_{t})\mid s_{0}=s\right]$, and the action-value function is $Q^{\pi}(s,a)=r(s,a)+\gamma\mathbb{E}_{s^{\prime}\sim P(\cdot|s,a)}[V^{\pi}(s^{\prime})]$. The optimal value function satisfies $V^{*}(s)=\max_{a\in\mathcal{A}}Q^{*}(s,a)$.
@@ 7.0.1 Adusumilli and Eckardt (2022): TD Learning for CCP Estimation
where $e(a,s)=\gamma_{\text{E}}-\ln P(a|s)$ under logit errors ($\gamma_{\text{E}}$ denoting the Euler constant), and the expectation is over the next-period state-action pair $(s^{\prime},a^{\prime})$ given the transition kernel $P(s^{\prime}|s,a)$ and the observed policy $P(a|s)$. Both $h$ and $g$ have a value-function structure; they satisfy a Bellman-like recursion under the observed (data-generating) policy, not the optimal policy.
This observation motivates the approach. Standard TD learning estimates value functions under a fixed policy from trajectory data, which is exactly what is needed here. Adusumilli et al. (2022) propose two methods.
The core idea is to approximate $h$ and $g$ with linear combinations of basis functions, then use regression on observed state-action sequences to estimate the weights. This reduces the infinite-dimensional fixed-point problem to a finite-dimensional linear system.
The first is the linear semi-gradient method. [^92] The method is called “semi-gradient” because it computes the gradient of only the prediction $\phi(a,s)^{\top}w$ with respect to $w$, not the full TD error including the bootstrap target $\phi(a^{\prime},s^{\prime})^{\top}w$. This avoids differentiating through the target but sacrifices guaranteed convergence in some settings. Approximate $h(a,s)\approx\phi(a,s)^{\top}w$ where $\phi:\mathcal{A}\times\mathcal{S}\to\mathbb{R}^{p}$ is a vector of basis functions (e.g., polynomials in state variables) and $w\in\mathbb{R}^{p}$ are the weights to be estimated. The TD(0) fixed-point equation is
@@
The sample analog replaces population expectations with averages over the observed panel:
@@
The second method is approximate value iteration (AVI). Rather than fixing a linear basis, AVI constructs a sequence of approximations $\hat{h}^{(k)}$ by iterating the Bellman-like operator and solving a nonparametric regression at each step. At iteration $k$, given the current approximation $\hat{h}^{(k-1)}$, one constructs pseudo-outcomes
@@
and then fits $\hat{h}^{(k)}$ by regressing $Y_{it}^{(k)}$ on $(a_{it},s_{it})$ using any machine learning method, including LASSO, random forests, or neural networks. This is the first DDC estimator compatible with arbitrary ML prediction methods, enabling application to very high-dimensional state spaces.
With $\hat{h}$ and $\hat{g}$ in hand, structural parameters are recovered by pseudo-maximum likelihood estimation (PMLE). For continuous state spaces, Adusumilli et al. (2022) derive a locally robust correction to the PMLE criterion that accounts for the nonparametric first-stage estimation of value terms, restoring $\sqrt{n}$-convergence of $\hat{\theta}$. [^93] An estimator has $\sqrt{n}$-convergence if its error shrinks at rate $n^{-1/2}$ where $n$ is the sample size. This is the standard parametric rate; slower rates (e.g., $n^{-1/4}$) indicate efficiency loss from nonparametric first stages. The PMLE score is $m(a,s;\theta,h,g)=\partial_{\theta}\ln\pi(a,s;\theta,h,g)$, where $\pi$ is the logit choice probability with continuation value $V(a,s)=h(a,s)^{\top}\theta+g(a,s)$. The naive estimator solves $\mathbb{E}_{n}[m(a,s;\theta,\hat{h},\hat{g})]=0$, but with continuous states this moment condition is not orthogonal to the first-stage estimates and converges slower than $\sqrt{n}$. The locally robust moment adds a debiasing correction:
@@
where $\lambda(a,s;\theta)$ solves a backward recursion that propagates the influence of estimation error through the dynamic structure. [^94] The term in braces is the temporal-difference error of the continuation value $V$. The adjoint $\lambda$ weights this TD error by its marginal impact on the PMLE score. See Online Appendix B.3 of Adusumilli et al. (2022) for the derivation. The corrected estimator $\hat{\theta}_{LR}$ solves $\mathbb{E}_{n}[\zeta_{n}]=0$ and is computationally no harder than the naive PMLE, since the correction is constant in $\theta$.
@@
Under regularity conditions, the linear semi-gradient estimator $\hat{h}$ satisfies $\|\hat{h}-h\|_{2}=O_{P}(n^{-1/2}(T-1)^{-1/2})$, where $\|\cdot\|_{2}$ denotes the $L^{2}(P)$ norm. [^95] The $L^{2}(P)$ norm is $\|f\|_{2}=(\int f(x)^{2}\,dP(x))^{1/2}$, measuring average squared deviation under the probability measure $P$. This is the natural norm for mean-squared-error analysis.
@@
The limitation is the linear semi-gradient’s dependence on the choice of basis functions, though cross-validation can guide this selection. AVI with flexible ML methods relaxes this constraint at the cost of slightly higher small-sample bias, which the locally robust correction addresses.
@@ 7.0.2 Hu and Yang (2025): Policy Gradient for DDC Estimation
Hu and Yang (2025) combine policy gradient methods with the Simulated Method of Moments (SMM) to estimate DDCs, with particular focus on models with unobserved state variables. Where the previous subsection replaced the inner fixed-point computation with TD learning, Hu and Yang replace it with policy gradient, sidestepping value function approximation entirely and working directly in the space of parametrized policies.
The outer loop is SMM. [^96] SMM (Gourieroux et al., 1993) estimates structural parameters by matching moments from simulated data to moments from observed data, avoiding direct evaluation of the likelihood function. Define a vector of data moments $\mathbf{M}_{d}$ computed from the observed panel. For candidate structural parameters $\theta$ and transition parameters $\xi$, simulate the model to produce simulated moments $\mathbf{M}_{s}(\theta,\xi)$. The estimator minimizes
@@
The inner loop parametrizes the choice probability directly as a logistic function of state variables. For a binary choice $J_{t}\in\{0,1\}$, the general form is $\Pr(J_{t}=1\mid\boldsymbol{X}_{t};\boldsymbol{\gamma}(\theta))=\text{logistic}(\boldsymbol{X}_{t}\boldsymbol{\gamma}(\theta))$, where $\boldsymbol{\gamma}(\theta)$ are policy parameters that depend on the structural parameters. [^97] The linear index can be replaced by higher-order terms of $\boldsymbol{X}_{t}$ or deep neural networks; the method requires only that the gradient $\nabla_{\boldsymbol{\gamma}}\log\pi_{\boldsymbol{\gamma}}$ has a closed form. In their application to a Rust bus engine model with unobserved bus condition $S_{t}^{*}$, this takes the form
@@
On an extended Rust bus engine model with a continuous unobserved bus condition following an AR(1) process, estimates of seven structural parameters are centered around their true values across 400 simulations (Table 1 in their paper). On a discrete-unobservable variant of the same model, the method matches the precision of Arcidiacono and Miller (2011)’s two-step EM algorithm, with computation time depending on the number of inner-loop iterations: at 5,000 iterations, RL+indirect inference runs in 0.40 minutes versus 0.66 minutes for CCP estimation; at 20,000 iterations, it rises to 1.15 minutes (Table 2 in their paper). Each outer-loop evaluation requires a full policy gradient convergence in the inner loop, so the computational advantage diminishes as more inner-loop iterations are used for precision.
@@ 7.1 Dynamic Oligopoly and Strategic Interaction
Dynamic oligopoly models combine elements of game theory and dynamic programming. Firms choose actions (prices, quantities, investment levels) strategically, anticipating both their own future states and the strategies of competitors. The state space grows combinatorially in the number of firms, and the requirement to compute best responses against other firms’ strategies adds further computational burden.

### 15 Mar 2026 v2

@@ 6 $\sim 10^{8}$ Failed
Where classical solutions exist, RL underperforms: the base-stock policy remains highly competitive when properly calibrated, and DRL required millions of training transitions to approach performance levels achievable through closed-form calculation. RL struggles particularly with multi-echelon coupling, where upstream orders affect downstream costs many periods later.127127127Credit assignment refers to the difficulty of determining which past actions caused a delayed reward or cost. In multi-echelon systems, an upstream ordering decision today may not affect customer-facing costs for many periods, making it difficult for RL to learn the causal connection. The value proposition for RL lies in problems where analytical solutions are unavailable: non-stationary demand, complex operational constraints, or cost structures that do not admit tractable decomposition. Even in these settings, successful deployment remains rare and requires extensive simulation infrastructure, domain expertise, and careful calibration against classical baselines.
@@ 8.6 Real-Time Bidding
Real-time bidding for display advertising presents a budget pacing problem: an advertiser must allocate a fixed budget $B_{0}$ across a campaign of $T$ auctions to maximize total conversions. Each impression is a second-price auction; the advertiser submits a bid and pays the second-highest bid if they win. The challenge is that bidding aggressively depletes budget early, while conservative bidding leaves value on the table.
Wu et al. (2018) formalized this as an MDP with state $s_{t}=(B_{t},t,w_{t})$, where $B_{t}$ is remaining budget, $t$ is auctions remaining, and $w_{t}$ is the recent win rate. The action is a bid multiplier $\lambda_{t}\in[\lambda_{\min},\lambda_{\max}]$, so the actual bid is $b_{t}=\lambda_{t}\cdot\bar{b}$, where $\bar{b}$ is a base bid calibrated to the estimated value-per-impression. The clearing price $c_{t}$ is drawn from a distribution $F_{c}$ estimated from historical data; the advertiser wins if $b_{t}\geq c_{t}$. The per-auction reward is:
$r_{t}=\mathbbm{1}\{b_{t}\geq c_{t}\}\cdot\mathrm{cvr}_{t},$ (68)
where $\mathrm{cvr}_{t}\in\{0,1\}$ is the conversion indicator. Budget evolves as $B_{t+1}=B_{t}-\mathbbm{1}\{b_{t}\geq c_{t}\}\cdot c_{t}$, and the episode terminates when $B_{t}\leq 0$ or t=0t=0.128128128The state space is discretized into budget bins and time bins. Wu et al. (2018) augment the state with bid landscape features derived from historical clearing price distributions, giving the agent information about current market competitiveness.
The agent is a deep Q-network trained on a simulator calibrated to production RTB data. Table 16 reports performance relative to rule-based pacing baselines. The DQN agent improves total conversions by conserving budget during high-competition periods and bidding aggressively when clearing prices are low, a pattern the rule-based approaches cannot learn.
Table 16: Real-time bidding results from Wu et al. (2018). Performance is relative to linear pacing on a simulator calibrated to production data.
Method Conversions vs. Linear Pacing Budget Utilization
Linear pacing baseline 98%
Hard threshold -8.3% 91%
Proportional (ECPC) ++4.1% 97%
DQN (Wu et al., 2018) ++11.7% 99%
@@ 8.7 Simulation Study: Bus Engine Replacement
I extend the single-engine problem to a fleet of $N$ engines with a capacity constraint limiting replacements per period. The state $s=(m_{1},\ldots,m_{N})$ records discretized mileage for each engine. Actions are subsets of engines to replace, subject to the capacity constraint. The per-period cost is $c(s,a)=\alpha\sum_{i}m_{i}+\beta|a|$, where $\alpha$ is the operating cost per unit mileage and $\beta$ is the replacement cost.129129129The mileage-dependent operating cost $\alpha\sum_{i}m_{i}$ follows Rust’s original specification $c(x,\theta_{1})=\theta_{11}x$, creating a non-trivial threshold replacement policy. The fleet extension uses deterministic mileage increments to isolate the combinatorial scaling challenge that arises from the joint state of multiple engines. Mileage evolves deterministically; replaced engines reset to $m=0$; others increment by one bin.130130130With $M=6$ mileage bins, the state space is $6^{N}$: 1,296 states at $N=4$, 7,776 at $N=5$, 46,656 at $N=6$. Value iteration is feasible for $N\leq 5$.
@@ 9 Structural Estimation with Reinforcement Learning
Several recent papers have used RL training loops, namely Q-learning, temporal-difference learning, policy gradient, and actor-critic methods, to solve structural economic models at scales where conventional dynamic programming fails.131131131I exclude papers that use neural network function approximation without an RL training mechanism. Inverse reinforcement learning is treated in the sister survey (Rust and Rawat, 2026). Throughout this chapter I adopt a unified notation. An MDP is a tuple $(\mathcal{S},\mathcal{A},P,r,\gamma)$ where $\mathcal{S}$ is the state space, $\mathcal{A}$ is the action space, $P(s^{\prime}|s,a)$ is the transition kernel, $r(s,a)$ is the per-period reward, and $\gamma\in[0,1)$ is the discount factor.132132132Several of the papers reviewed here use $\beta$ for the discount factor, following economics convention. I translate all results to $\gamma$ for consistency with the RL literature and the rest of this survey. A policy $\pi:\mathcal{S}\to\Delta(\mathcal{A})$ maps states to distributions over actions. The value function under $\pi$ is $V^{\pi}(s)=\mathbb{E}_{\pi}\left[\sum_{t=0}^{\infty}\gamma^{t}r(s_{t},a_{t})\mid s_{0}=s\right]$, and the action-value function is $Q^{\pi}(s,a)=r(s,a)+\gamma\mathbb{E}_{s^{\prime}\sim P(\cdot|s,a)}[V^{\pi}(s^{\prime})]$. The optimal value function satisfies $V^{*}(s)=\max_{a\in\mathcal{A}}Q^{*}(s,a)$.
The canonical structural estimation framework for MDPs was formulated by Rust (1994), whose nested fixed-point (NFXP) algorithm embeds the Bellman equation inside a maximum likelihood estimator. The methods reviewed in this chapter replace the inner fixed-point computation with RL-based approximations.
Figure 16: NFXP versus RL-based structural estimation. Top: the nested fixed-point algorithm evaluates the likelihood by solving the Bellman equation to convergence inside each optimizer step. Bottom: single-loop stochastic approximation updates structural parameters $\theta$ and value/policy weights $\omega$ simultaneously from data batches. The NFXP algorithm is due to Rust (1987).
@@ 9.0.1 Adusumilli and Eckardt (2022): TD Learning for CCP Estimation
where $e(a,s)=\gamma_{\text{E}}-\ln P(a|s)$ under logit errors ($\gamma_{\text{E}}$ denoting the Euler constant), and the expectation is over the next-period state-action pair $(s^{\prime},a^{\prime})$ given the transition kernel $P(s^{\prime}|s,a)$ and the observed policy $P(a|s)$. Both $h$ and $g$ satisfy a Bellman-like recursion under the observed (data-generating) policy, not the optimal policy, so standard TD learning applies directly. Adusumilli et al. (2022) propose two methods.
The first is the linear semi-gradient method.133133133The method is called “semi-gradient” because it computes the gradient of only the prediction $\phi(a,s)^{\top}w$ with respect to $w$, not the full TD error including the bootstrap target $\phi(a^{\prime},s^{\prime})^{\top}w$. This avoids differentiating through the target but sacrifices guaranteed convergence in some settings. Approximate $h(a,s)\approx\phi(a,s)^{\top}w$ where $\phi:\mathcal{A}\times\mathcal{S}\to\mathbb{R}^{p}$ is a vector of basis functions (e.g., polynomials in state variables) and $w\in\mathbb{R}^{p}$ are the weights to be estimated. The TD(0) fixed-point equation is
@@
The sample analog replaces population expectations with averages over the observed panel.
@@
The second method is approximate value iteration (AVI), which iterates the Bellman-like operator using nonparametric regression. At iteration $k$, one constructs pseudo-outcomes
@@
and then fits $\hat{h}^{(k)}$ by regressing $Y_{it}^{(k)}$ on $(a_{it},s_{it})$ using any machine learning method, including LASSO, random forests, or neural networks.134134134LASSO (Tibshirani, 1996) adds an $\ell_{1}$ penalty to a regression loss, shrinking many coefficients to zero for sparse solutions; random forests average many decision trees fit to random subsamples and feature subsets for nonparametric estimation; neural networks compose affine transformations with elementwise nonlinearities across multiple layers. This is the first DDC estimator compatible with arbitrary ML prediction methods, enabling application to very high-dimensional state spaces.
With $\hat{h}$ and $\hat{g}$ in hand, structural parameters are recovered by pseudo-maximum likelihood estimation (PMLE). For continuous state spaces, Adusumilli et al. (2022) derive a locally robust correction to the PMLE criterion that accounts for the nonparametric first-stage estimation of value terms, restoring $\sqrt{n}$-convergence of $\hat{\theta}$.135135135An estimator has $\sqrt{n}$-convergence if its error shrinks at rate $n^{-1/2}$ where $n$ is the sample size. This is the standard parametric rate; slower rates (e.g., $n^{-1/4}$) indicate efficiency loss from nonparametric first stages. The PMLE score is $m(a,s;\theta,h,g)=\partial_{\theta}\ln\pi(a,s;\theta,h,g)$, where $\pi$ is the logit choice probability with continuation value $V(a,s)=h(a,s)^{\top}\theta+g(a,s)$. The naive estimator solves $\mathbb{E}_{n}[m(a,s;\theta,\hat{h},\hat{g})]=0$, but with continuous states this moment condition is not orthogonal to the first-stage estimates and converges slower than $\sqrt{n}$. The locally robust moment adds a debiasing correction:
@@
where $\lambda(a,s;\theta)$ solves a backward recursion that propagates the influence of estimation error through the dynamic structure.136136136The term in braces is the temporal-difference error of the continuation value $V$. The adjoint $\lambda$ weights this TD error by its marginal impact on the PMLE score. See Online Appendix B.3 of Adusumilli et al. (2022) for the derivation. The corrected estimator $\hat{\theta}_{LR}$ solves $\mathbb{E}_{n}[\zeta_{n}]=0$ and is computationally no harder than the naive PMLE, since the correction is constant in $\theta$.
@@
Under regularity conditions, the linear semi-gradient estimator $\hat{h}$ satisfies $\|\hat{h}-h\|_{2}=O_{P}(n^{-1/2}(T-1)^{-1/2})$, where $\|\cdot\|_{2}$ denotes the $L^{2}(P)$ norm.137137137The $L^{2}(P)$ norm is $\|f\|_{2}=(\int f(x)^{2}\,dP(x))^{1/2}$, measuring average squared deviation under the probability measure $P$. This is the natural norm for mean-squared-error analysis.
@@ 9.0.2 Hu and Yang (2025): Policy Gradient for DDC Estimation
Hu and Yang (2025) combine policy gradient methods with the Simulated Method of Moments (SMM) to estimate DDCs, with particular focus on models with unobserved state variables.
The outer loop is SMM.138138138SMM (Gourieroux et al., 1993) estimates structural parameters by matching moments from simulated data to moments from observed data, avoiding direct evaluation of the likelihood function. Define a vector of data moments $\mathbf{M}_{d}$ computed from the observed panel. For candidate structural parameters $\theta$ and transition parameters $\xi$, simulate the model to produce simulated moments $\mathbf{M}_{s}(\theta,\xi)$. The estimator minimizes
@@
The inner loop parametrizes the choice probability directly as a logistic function of state variables. For a binary choice $J_{t}\in\{0,1\}$, the general form is $\Pr(J_{t}=1\mid\boldsymbol{X}_{t};\boldsymbol{\gamma}(\theta))=\text{logistic}(\boldsymbol{X}_{t}\boldsymbol{\gamma}(\theta))$, where $\boldsymbol{\gamma}(\theta)$ are policy parameters that depend on the structural parameters.139139139The linear index can be replaced by higher-order terms of $\boldsymbol{X}_{t}$ or deep neural networks; the method requires only that the gradient $\nabla_{\boldsymbol{\gamma}}\log\pi_{\boldsymbol{\gamma}}$ has a closed form. In their application to a Rust bus engine model with unobserved bus condition $S_{t}^{*}$, this takes the form
@@
On an extended Rust bus engine model with a continuous unobserved bus condition following an AR(1) process, estimates of seven structural parameters are centered around their true values across 400 simulations. On a discrete-unobservable variant, the method matches the precision of Arcidiacono and Miller (2011)’s two-step EM algorithm at comparable computation times, though the advantage diminishes as more inner-loop iterations are used for precision.
@@ 9.1 Dynamic Oligopoly and Strategic Interaction
Dynamic oligopoly models combine game theory and dynamic programming: firms choose actions strategically while anticipating competitors’ strategies, and the state space grows combinatorially in the number of firms.